OpenStudy (dls):

Which of the following cannot be the resultant of the vectors of magnitude 5 and 10?

OpenStudy (dls):

@experimentX @mukushla @Yahoo! help help!

OpenStudy (dls):

7 5 8 2

OpenStudy (anonymous):

we must use an inequality for \(|\vec a+\vec b|\)

OpenStudy (dls):

Didn't get it :/

OpenStudy (anonymous):

also we can prove \[|\vec a|-|\vec b| \le|\vec a+\vec b| \le |\vec a|+|\vec b|\]

OpenStudy (dls):

so what should be your answer?

OpenStudy (anonymous):

i think it is 2

OpenStudy (anonymous):

Do u have the answer key

OpenStudy (dls):

no..

OpenStudy (anonymous):

\[|\vec a|-|\vec b| \le|\vec a+\vec b| \le |\vec a|+|\vec b|\]\[10-5 \le|\vec a+\vec b| \le 10+5\]\[5 \le|\vec a+\vec b| \le 15\]so it can not be 2

OpenStudy (anonymous):

just look this the resultant should not be less than the given two vectors

OpenStudy (anonymous):

but better to go with Expers way to understand the inequality

OpenStudy (dls):

where did we get that inequality from?